Klapperich Lab:Notebook/Lab Meeting Notes/2008/11/04

From OpenWetWare
Jump to navigationJump to search
Owwnotebook icon.png Project name Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

4 November 2008 Lab Meeting Agenda

† Mark's research report today.

† Announcements

  • 720 Migration in full swing. Make sure things you do not want "shared" are clearly marked. Check with Sonali before using any space in 709.
  • ALL - please heed lab safety rules. Re-read Sonali's email of earlier this week!

† Flu R01

  • No amendment needed. Flu meeting on 10/31 to discuss new procedures.
  • Plaque assay status?
  • Cathie Submitted the progress report to INSPIR.

† SEPSIS Project - CIMIT Bird Flu

  • Run E. coli RT-PCR using primers for a genomic target.
  • Report DNA amounts along with the RT-PCR data for the Sepsis paper.

†RNA project

  • Jeff Braman will visit in early Dec.
  • Do dilutions. Send to Bioanalyzer.


  • Mark finish control circuit
  • Bacteria Experiment. Failed first time. Any progress?
  • Jane, Cathie, Larry, Ranjith met again. JZ will try process with teflon substrates.

† Valve Array/Valve building (FJ)

  • Waiting for fixture still from Melissa. Hopefully this week. UPDATE?
  • Received schematic and quote. Frank met with Dr. Do to go over schematic and see what's next. Met with Sonali (3 Nov) to better delineate system requirements, will meet with Mark to coordinate fixture/pressure-driven flow system interfacing before submitting revised specs to Superior Fluid Solutions.

† Fraunhofer: LOAC

  • Update?
  • Alex working on more toluene bonding issues.
  • Try Ticona material. (HM, JD, AG)
  • Still Still Still troubleshooting ??

† Biointerfaces group

  • Mask writer's calibration was done.
  • Paul Mak told that Mask aligner at Photonics center has UV filter. Try that machine.
  • Preliminary E. coli experiment is ongoing.
  • Try DAPI bacteria stain to visualize e-coli soon.

† CIMIT- Colson Grant

  • Preliminary plan for IRB submission at Harvard in place. Target date 1/1/09
  • STAR-CD Renewal - January ($2K)


  • QQ gel issues. Resolved??
  • Only 1/5 chips are good at the hot embossing step. The vaccum chuck will be checked and o-rings changed - Will try after fixes and report next week. UPDATE?

* Fixuture for PCR chip. The heater and the thermocouples would be on the fixture.

  • Qingqing has to do experiments with MinCheol's designs. AS SOON AS CLEANING CONTEROLS ARE IN PLACE.
  • Mincheol will work on DNA migration simulation after IMECE 2008.


  • Cathie submitted One pager Whitepaper. Waiting for response.

† Senior project

  • Megan ordering oligos and primers for the RCA. Quote. Check with Heiko on his sources.
  • Haven't seen Teddy yet. What is up with him?

† F31: Cochlea

  • Received paper edits from Doug.
  • Make changes and give to Cathie ASAP.
  • Journal chosen?
  • Need to order Alpha-Tectorin protein.

† Silica Optimization (Lambda):

  • Ongoing. More concentrations. .
  • Pressure testing on these channels when they are flowing.